OpenStudy (anonymous):

express in simplest form with a prime number base 2^t*8

OpenStudy (unklerhaukus):

is that \(2^{t\times8}\) or \(2^{t}\times8\) ?

OpenStudy (unklerhaukus):

if its \(2^t\times8\). use \[\large x^n\times x^m=x^{n+m}\]

OpenStudy (unklerhaukus):

and 8=2×2×2